Tohban Report 2011-05-18

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 11 May 2011
End Date: 18 May 2011
Tohban: Claire Raftery
Tohban email:
Next Tohban: Jim Mc Tiernan
List all reports


Solar Activity

Activity was quiet this week despite having up to 6 active regions on disk at any one time - C class activity was the highlight of the week. Stereo B shows this level of activity is likely to continue, with some moderate sized ARs set to rotate onto disk in the coming days.

How many GOES flares occurred?

Flares above B, C, M, X class were     20     3     0     0

And how many of these are listed in the RHESSI flare list?

Flares above B, C, M, X class were     10     3     0     0

And how many had EXCELLENT coverage?

Flares above B, C, M, X class were      7     0     0     0

There were RHESSI flares/GOES flares 50 / 23 over the time range 11-May-11 18-May-11

Memory Management

The SSR is emptying at the end of each pass. We remain in normal/active decimation mode.

Data Gaps

No significant data gaps this week.

GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2011-05-14T21:15:00.000 -- 2011-05-14T21:20:00.000 300.00000 NO GAPS IN APP_ID = 154 (VC1-PMTRAS) NO GAPS IN APP_ID = 102 (VC3-MONITOR RATES)

Spacecraft Status

Arcs have been appearing in Detector 6 again, sometimes sparked by a spike eg: but not always. The high voltage was lowered by four counts on Monday but this was insufficient. As such, it was lowered by a further 6 steps today.

Personal tools